OpenStudy (anonymous):

The rate of growth of a particular population is given by dP/dt = 50t^2 - 100t^(3/2) , where P is population size and t is the time in years. Assume the initial population is 25,000. Determine the population function, P(t).

OpenStudy (kagıtucak):

I cannot see the equation after the phrase "given by"

OpenStudy (anonymous):

sorry about that!

OpenStudy (kagıtucak):

You can take the integral of dP/dt to get the function P(t). Lets do it: P(t)=(50/3)*(t^3/3)-(100*2/5)*(t^(5/2))+c P(t)=(50/3)*(t^3/3)-(40)*(t^(5/2))+c We are given that initial population is 25000. Lets put it in t=0: 25000=0-0+c c=25000 Hence, P(t)=(50/3)*(t^3/3)-(40)*(t^(5/2))+25000 Hope I didnt make a mistake:)